# Robust Bayesian Scene Reconstruction by Leveraging Retrieval-Augmented Priors

### v2

@@ VII Conclusion
We introduced the concept of retrieval-augmented priors, where we retrieve relevant components of a prior distribution during Bayesian inference. We also introduced a novel Bayesian method for scene reconstruction that uses informative priors. Our method, BRRP, leveraged existing mesh datasets to build its prior and probabilistically reconstructed the scene leveraging these priors in a retrieval-augmented manner. We experimentally showed our method was more robust than a deep learning method as well as more accurate than a mapping method. Finally, we showed a qualitative example of recovering principled uncertainty from our method.

### v3

@@ VII Conclusion
We introduced the concept of retrieval-augmented priors, where we retrieve relevant components of a prior distribution during Bayesian inference. We also introduced a novel Bayesian method for scene reconstruction that uses informative priors. Our method, BRRP, leveraged existing mesh datasets to build its prior and probabilistically reconstructed the scene leveraging these priors in a retrieval-augmented manner. We showed our method was more robust than a deep learning method as well as more accurate than a mapping method. We also showed a qualitative example of recovering principled uncertainty from our method. Finally, we demonstrated that our robust reconstructions improve downstream performance, where BRRP reconstructions enable precise collision avoidance as the robot moves to execute grasps.